# Robust Molecular Property Prediction via Densifying Scarce Labeled Data

### v1

@@ Abstract
A widely recognized limitation of molecular prediction models is their reliance on structures observed in the training data, resulting in poor generalization to out-of-distribution compounds. Yet in drug discovery, the compounds most critical for advancing research often lie beyond the training set, making the bias toward the training data particularly problematic. This mismatch introduces substantial covariate shift, under which standard deep learning models produce unstable and inaccurate predictions. Furthermore, the scarcity of labeled data—stemming from the onerous and costly nature of experimental validation—further exacerbates the difficulty of achieving reliable generalization. To address these limitations, we propose a novel meta-learning-based approach that leverages unlabeled data to interpolate between in-distribution (ID) and out-of-distribution (OOD) data, enabling the model to meta-learn how to generalize beyond the training distribution. We demonstrate significant performance gains over state-of-the-art methods on challenging real-world datasets that exhibit substantial covariate shift.
@@
Figure 1: Concept. We densify the train dataset using external unlabeled data (context point) for robust generalization across covariate shift. Further information on the notation is available in Section 4.
@@ 1 Introduction
Molecular property prediction plays a central role in drug discovery pipelines, enabling researchers to prioritize compounds for costly and time-consuming experimental validation. Accurate computational models have the potential to dramatically accelerate early-stage discovery by predicting critical attributes such as bioactivity, toxicity, and solubility before synthesis (Schneider, 2018; Vamathevan et al., 2019). However, building reliable predictive models that generalize to novel, unseen compounds remains a fundamental challenge.
Standard molecular property prediction models tend to rely heavily on patterns observed within the training distribution, resulting in poor generalization to out-of-distribution compounds (Klarner et al., 2023; Ovadia et al., 2019; Koh et al., 2021). In drug discovery, this limitation is particularly problematic, since the compounds most crucial for advancing research often lie far beyond the chemical spaces represented during training (Lee et al., 2023). The resulting covariate shift introduces significant obstacles to reliable prediction, with models frequently producing unstable outputs when extrapolating to new regions of chemical space. Further compounding these challenges, experimental validation of molecular properties is both costly and resource-intensive, leading to a scarcity of labeled data and increasing reliance on computational exploration (Altae-Tran et al., 2017). In addition, available labeled data is typically concentrated in narrow regions of chemical space, introducing bias that hampers generalization to unseen compounds (Klarner et al., 2023).
While vast collections of unlabeled molecular structures are readily available (Sterling & Irwin, 2015; Kim et al., 2021), offering rich information about the structure of chemical space, existing methods often fail to fully exploit this resource to improve generalization under covariate shift (Klarner et al., 2023). Therefore, we propose a novel meta-learning based method that leverages unlabeled data to densify the scarce train dataset and guide the model toward sensible behavior in unexplored regions of chemical space.
@@ 2 Methodology
Model HIVPROT DPP4 NK1 count vector bit vector count vector bit vector count vector bit vector $L_{1}$-Regression $1.137{\scriptstyle\pm.000}$ $0.714{\scriptstyle\pm.000}$ $1$000 $1.130{\scriptstyle\pm.000}$ $0.482{\scriptstyle\pm.000}$ $0.442{\scriptstyle\pm.000}$ $L_{2}$-Regression 0.999±.000plus-or-minus0.999.0000.$9$000 $0.723{\scriptstyle\pm.000}$ $1.495{\scriptstyle\pm.000}$ $1.143{\scriptstyle\pm.000}$ $0.498{\scriptstyle\pm.000}$ $0.436{\scriptstyle\pm.000}$ Random Forest $0.815{\scriptstyle\pm.009}$ $0.834{\scriptstyle\pm.010}$ $1.473{\scriptstyle\pm.008}$ $1.461{\scriptstyle\pm.012}$ $0.458{\scriptstyle\pm.002}$ $0.438{\scriptstyle\pm.002}$ MLP $0.768{\scriptstyle\pm.014}$ $2.118{\scriptstyle\pm.015}$ $1.393{\scriptstyle\pm.024}$ $1.094{\scriptstyle\pm.029}$ $0.443{\scriptstyle\pm.007}$ $0.399{\scriptstyle\pm.006}$ Mixup $0.764{\scriptstyle\pm 0.008}$ $0.691{\scriptstyle\pm 0.022}$ $1.439{\scriptstyle\pm 0.021}$ $1.212{\scriptstyle\pm 0.012}$ $0.481{\scriptstyle\pm 0.002}$ $0.479{\scriptstyle\pm 0.003}$ Mixup (w/ outlier exposure) $0.748{\scriptstyle\pm 0.01}$ $0.677{\scriptstyle\pm 0.015}$ $1.384{\scriptstyle\pm 0.012}$ $1.224{\scriptstyle\pm 0.016}$ $0.442{\scriptstyle\pm 0.005}$ $0.443{\scriptstyle\pm 0.005}$ Manifold Mixup $0.88{\scriptstyle\pm 0.023}$ $0.898{\scriptstyle\pm 0.022}$ $1.414{\scriptstyle\pm 0.021}$ $1.367{\scriptstyle\pm 0.04}$ $0.432{\scriptstyle\pm 0.005}$ $0.499{\scriptstyle\pm 0.013}$ Manifold Mixup (w/ bilevel) $0.484{\scriptstyle\pm 0.011}$ $0.959{\scriptstyle\pm 0.257}$ $1.19{\scriptstyle\pm 0.05}$ $1.217{\scriptstyle\pm 0.068}$ $0.43{\scriptstyle\pm 0.013}$ $0.536{\scriptstyle\pm 0.039}$ Q-SAVI $0.682{\scriptstyle\pm.019}$ $0.664{\scriptstyle\pm.028}$ $1.332{\scriptstyle\pm.017}$ $1.028{\scriptstyle\pm.027}$ $0.436{\scriptstyle\pm.007}$ $0.387{\scriptstyle\pm.012}$ Ours (Deepsets) $0.555\pm 0.096$ $\mathbf{0.364\pm 0.018}$ $\mathbf{0.984\pm 0.018}$ $\mathbf{0.963\pm 0.017}$ $0.455\pm 0.016$ $\mathbf{0.376\pm 0.008}$ Ours (Set Trans.) $\mathbf{0.39\pm 0.011}$ $0.726\pm 0.159$ $1.121\pm 0.037$ $0.986\pm 0.021$ $\mathbf{0.429\pm 0.01}$ $0.397\pm 0.015$
We utilize bilevel optimization for training meta-learner $f_{\theta_{l}}$, and treat the set function parameter $\mu_{\lambda}$ as a hyperparameter to be optimized in the outer loop (Lorraine et al., 2019). As shown in Table 2 (w/o bilevel optimization), simply optimizing the meta-learner parameters $\theta$ and the set function parameters $\lambda$ jointly can lead to overfitting to the task distribution and harms test-time generalization. Following the setting of (Lorraine et al., 2019), during training, we only update the parameter $\theta$ in the inner loop and we only update the parameter $\lambda$ in the outer loop.
@@
where $K$ is a hyperparameter of the samples drawn from $\mathcal{D}_{\text{unlabeled}}$ for each minibatch. In Table 2, we empirically show that using a random $y_{i,k}^{(\text{mvalid})}\sim\mathcal{N}(0,1)$ achieves better or comparable performance to using a labeled $y_{i,k}^{(\text{mvalid})}\sim\mathcal{D}_{\text{oracle}}$ corresponding to the real label for $x_{i,k}^{(\text{mvalid})}$ from the unlabeled dataset which has been labeled by an oracle. This setting aligns with the role of $L_{V}$ in guiding the set function $\mu_{\lambda}$ to generalize to out-of-distribution data, as sampling pseudo-labels from $\mathcal{N}(0,1)$ introduces controlled label noise that regularizes the model. The outer loop loss $L_{V}$ can then be expressed as a function of two variables $L_{V}(\lambda,\theta^{*}(\lambda))$, and therefore, final contribution of the hypergradient is given by,
@@
Luckily, the inverse Hessian vector product can be approximated and computed efficiently using automatic differentiation and Neuman series iterations (Lorraine et al., 2019). In Section 4, we visually display how the model $\hat{f}_{\theta,\lambda}$ effectively uses $\mathcal{D}_{\text{context}}$ to densify the data distribution around $\mathcal{D}_{\text{train}}$.
As the mixing of $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{context}}$ is meant to densify the sparse training data, we only perform the mixing at train time in the inner loop. At test time and in the outer loop optimization, the set function $\mu_{\lambda}$ receives an input as a singleton set and performs no mixing with any exogeneous distribution.
@@ 3 Experiments
Model w/ ctx. w/ bilevel $y^{(\text{mvalid})}$ HIVPROT DPP4 NK1 count vector bit vector count vector bit vector count vector bit vector Ours (Deepsets) rand $1.147\pm 0.055$ $1.012\pm 0.025$ $1.342\pm 0.019$ $1.344\pm 0.039$ $0.45\pm 0.01$ $0.494\pm 0.006$ Ours (Set Trans.) rand $1.193\pm 0.038$ $1.008\pm 0.03$ $1.34\pm 0.026$ $1.359\pm 0.013$ $0.482\pm 0.002$ $0.487\pm 0.009$ Ours (MLP) rand $1.043\pm 0.013$ $1.049\pm 0.013$ $1.42\pm 0.009$ $1.269\pm 0.017$ $0.435\pm 0.002$ $0.466\pm 0.004$ Ours (Deepsets) real $0.517\pm 0.053$ $0.402\pm 0.028$ $\mathbf{0.977\pm 0.018}$ $\mathbf{0.942\pm 0.023}$ $0.505\pm 0.045$ $0.412\pm 0.021$ Ours (Set Trans.) real $0.55\pm 0.083$ $0.52\pm 0.046$ $0.996\pm 0.016$ $0.969\pm 0.025$ $0.624\pm 0.072$ $0.711\pm 0.071$ Ours (Deepsets) rand $0.555\pm 0.096$ $\mathbf{0.364\pm 0.018}$ $0.984\pm 0.018$ $0.963\pm 0.017$ $0.455\pm 0.016$ $\mathbf{0.376\pm 0.008}$ Ours (Set Trans.) rand $\mathbf{0.39\pm 0.011}$ $0.726\pm 0.159$ $1.121\pm 0.037$ $0.986\pm 0.021$ $\mathbf{0.429\pm 0.01}$ $0.397\pm 0.015$
@@
In addition to the baselines presented in (Klarner et al., 2023), to empirically show the capability of our set function $\mu_{\lambda}$, we evaluate two related state-of-the-art interpolation methods—Mixup (Zhang et al., 2017) and Manifold Mixup (Verma et al., 2019). As mixup performs the mixing procedure in the input space, there are no parameters to train for bilevel optimization. We therefore report two variants of mixup (plain and outlier exposure) where outlier exposure is exposed to the same $\mathcal{D}_{\text{mvalid}}$ with random labels as described above. Manifold mixup performs the mixup procedure in the feature space between layers, therefore we can straightforwardly apply our bilevel optimization procedure to manifold mixup. Therefore, we report two variants of manifold mixup (plain and bilevel optimization).
@@ 4 Analysis
To visually assess how our model leverages $\mathcal{D}_{\text{context}}$ to densify $\mathcal{D}_{\text{train}}$, and to compare model behaviors on out-of-distribution (OOD) inputs, we visualize the penultimate-layer embeddings $\mathbf{z}_{i}=x_{i}^{(L-1)}$ , using t-SNE (Van der Maaten & Hinton, 2008) across three models that interpolate $\mathcal{D}_{\text{train}}$ with $\mathcal{D}_{\text{context}}$: Ours, Mixup (w/ outlier exposure), and Manifold Mixup (w/ bilevel optim.), all trained on the DPP4 bit-vector dataset. We randomly sample $x_{i}\sim\mathcal{D}_{\text{train}}$, $C_{i}=\{c_{ij}\}_{j=1}^{100}\sim\mathcal{D}_{\text{context}}$, $x_{i}^{\text{ood}}\sim\mathcal{D}_{\text{ood}}$, where $\mathcal{D}_{\text{ood}}$ signifies an out-of-distribution dataset which was never seen during training. In Figure 2, each color indicates a penultimate-layer embedding $\mathbf{z}\in\mathbf{Z}$ under different input configurations. We denote the embeddings as follows: $\mathbf{Z}_{\text{joint}}$ for orange ($\{x_{i},C_{i}\}$), $\mathbf{Z}_{\text{input}}$ for blue ($x_{i}$ alone), $\mathbf{Z}_{\text{context}}$ for green ($C_{i}$ alone), and $\mathbf{Z}_{\text{ood}}$ for purple (OOD samples $x_{i}^{\text{ood}}$). More t-SNE visualizations and further details are provided in Appendix B.

### v2

@@ Abstract
A widely recognized limitation of molecular prediction models is their reliance on structures observed in the training data, resulting in poor generalization to out-of-distribution compounds. Yet in drug discovery, the compounds most critical for advancing research often lie beyond the training set, making the bias toward the training data particularly problematic. This mismatch introduces substantial covariate shift, under which standard deep learning models produce unstable and inaccurate predictions. Furthermore, the scarcity of labeled data—stemming from the onerous and costly nature of experimental validation—further exacerbates the difficulty of achieving reliable generalization. To address these limitations, we propose a novel meta-learning-based approach that leverages unlabeled data to interpolate between in-distribution (ID) and out-of-distribution (OOD) data, enabling the model to meta-learn how to generalize beyond the training distribution. We demonstrate significant performance gains on challenging real-world datasets with substantial covariate shift, supported by t-SNE visualizations highlighting our interpolation method.
@@
Figure 1: Concept. We densify the train dataset using external unlabeled data (context point) for robust generalization across covariate shift. Notation details are provided in Section 4.
@@ 1 Introduction
Molecular property prediction plays a central role in drug discovery pipelines, enabling researchers to prioritize compounds for costly and time-consuming experimental validation. Accurate computational models have the potential to dramatically accelerate early-stage discovery by predicting critical attributes such as bioactivity, toxicity, and solubility before synthesis (Schneider, 2018; Vamathevan et al., 2019). However, building reliable predictive models generalizing to novel, unseen compounds remains a fundamental challenge.
Standard molecular property prediction models tend to rely heavily on patterns observed within the training distribution, resulting in poor generalization to out-of-distribution compounds (Klarner et al., 2023; Ovadia et al., 2019; Koh et al., 2021). In drug discovery, this limitation is particularly problematic, since the compounds most crucial for advancing research often lie far beyond the chemical spaces represented during training (Lee et al., 2023). The resulting covariate shift introduces significant obstacles to reliable prediction, with models frequently producing unstable outputs when extrapolating to new regions of chemical space. Further compounding these challenges, experimental validation of molecular properties is both costly and resource-intensive, leading to a scarcity of labeled data and increasing reliance on computational exploration (Altae-Tran et al., 2017). Also, available labeled data is typically concentrated in narrow regions of chemical space, introducing bias that hampers generalization to unseen compounds (Klarner et al., 2023).
While vast collections of unlabeled molecular structures are readily available (Sterling & Irwin, 2015; Kim et al., 2021), offering rich information about the structure of chemical space, existing methods often fail to fully exploit this resource to improve generalization (Klarner et al., 2023). Therefore, we propose a novel meta-learning based method that leverages unlabeled data to densify the scarce train dataset and guide the model toward sensible behavior in unexplored regions of chemical space. Our code can be found at https://github.com/JinA0218/drugood-densify.
@@ 2 Methodology
Model HIVPROT DPP4 NK1 count vector bit vector count vector bit vector count vector bit vector $L_{1}$-Regression $1.137{\scriptstyle\pm.000}$ $0.714{\scriptstyle\pm.000}$ $1$000 $1.130{\scriptstyle\pm.000}$ $0.482{\scriptstyle\pm.000}$ $0.442{\scriptstyle\pm.000}$ $L_{2}$-Regression 0.999±.000plus-or-minus0.999.0000.$9$000 $0.723{\scriptstyle\pm.000}$ $1.495{\scriptstyle\pm.000}$ $1.143{\scriptstyle\pm.000}$ $0.498{\scriptstyle\pm.000}$ $0.436{\scriptstyle\pm.000}$ Random Forest $0.815{\scriptstyle\pm.009}$ $0.834{\scriptstyle\pm.010}$ $1.473{\scriptstyle\pm.008}$ $1.461{\scriptstyle\pm.012}$ $0.458{\scriptstyle\pm.002}$ $0.438{\scriptstyle\pm.002}$ MLP $0.768{\scriptstyle\pm.014}$ $2.118{\scriptstyle\pm.015}$ $1.393{\scriptstyle\pm.024}$ $1.094{\scriptstyle\pm.029}$ $0.443{\scriptstyle\pm.007}$ $0.399{\scriptstyle\pm.006}$ Mixup $0.764{\scriptstyle\pm 0.008}$ $0.691{\scriptstyle\pm 0.022}$ $1.439{\scriptstyle\pm 0.021}$ $1.212{\scriptstyle\pm 0.012}$ $0.481{\scriptstyle\pm 0.002}$ $0.479{\scriptstyle\pm 0.003}$ Mixup (w/ outlier exposure) $0.748{\scriptstyle\pm 0.01}$ $0.677{\scriptstyle\pm 0.015}$ $1.384{\scriptstyle\pm 0.012}$ $1.224{\scriptstyle\pm 0.016}$ $0.442{\scriptstyle\pm 0.005}$ $0.443{\scriptstyle\pm 0.005}$ Manifold Mixup $0.88{\scriptstyle\pm 0.023}$ $0.898{\scriptstyle\pm 0.022}$ $1.414{\scriptstyle\pm 0.021}$ $1.367{\scriptstyle\pm 0.04}$ $0.432{\scriptstyle\pm 0.005}$ $0.499{\scriptstyle\pm 0.013}$ Manifold Mixup (w/ bilevel) $0.484{\scriptstyle\pm 0.011}$ $0.804{\scriptstyle\pm 0.086}$ $1.19{\scriptstyle\pm 0.05}$ $1.217{\scriptstyle\pm 0.068}$ $0.43{\scriptstyle\pm 0.013}$ $0.536{\scriptstyle\pm 0.039}$ Q-SAVI $0.682{\scriptstyle\pm.019}$ $0.664{\scriptstyle\pm.028}$ $1.332{\scriptstyle\pm.017}$ $1.028{\scriptstyle\pm.027}$ $0.436{\scriptstyle\pm.007}$ $0.387{\scriptstyle\pm.012}$ Ours (Deepsets) $0.555\pm 0.096$ $\mathbf{0.364\pm 0.018}$ $\mathbf{0.984\pm 0.018}$ $\mathbf{0.963\pm 0.017}$ $0.455\pm 0.016$ $\mathbf{0.376\pm 0.008}$ Ours (Set Trans.) $\mathbf{0.39\pm 0.011}$ $0.726\pm 0.159$ $1.121\pm 0.037$ $0.986\pm 0.021$ $\mathbf{0.429\pm 0.01}$ $0.397\pm 0.015$
We utilize bilevel optimization for training meta-learner $f_{\theta_{l}}$, and treat the set function parameter $\mu_{\lambda}$ as a hyperparameter to be optimized in the outer loop (Lorraine et al., 2019). As shown in Table 2 (w/o bilevel optimization), simply optimizing the meta-learner parameters $\theta$ and the set function parameters $\lambda$ jointly can lead to overfitting to the task distribution and harms test-time generalization. Following the setting of (Lorraine et al., 2019), during training, we only update the parameter $\theta$ in the inner loop and we only update the parameter $\lambda$ in the outer loop (see Figure 9(b) for the detailed model structure of the bilevel optimization).
@@
where $K$ is a hyperparameter of the samples drawn from $\mathcal{D}_{\text{unlabeled}}$ for each minibatch. In Table 2, we empirically show that using a random $y_{i,k}^{(\text{mvalid})}\sim\mathcal{N}(0,1)$ achieves better or comparable performance to using a labeled $y_{i,k}^{(\text{mvalid})}\sim\mathcal{D}_{\text{oracle}}$ corresponding to the real label for $x_{i,k}^{(\text{mvalid})}$ from the unlabeled dataset which has been labeled by an oracle. This setting aligns with the role of $L_{V}$ in guiding the set function $\mu_{\lambda}$ to generalize to out-of-distribution data, as sampling pseudo-labels from $\mathcal{N}(0,1)$ introduces controlled label noise that regularizes the model. The outer loop loss $L_{V}$ can then be expressed as a function of two variables $L_{V}(\lambda,\theta^{*}(\lambda))$, and therefore, the hypergradient is given by,
@@
Luckily, the inverse Hessian vector product can be approximated and computed efficiently using automatic differentiation and Neuman series iterations (Lorraine et al., 2019). In Section 4, we display how the model $\hat{f}_{\theta,\lambda}$ effectively uses $\mathcal{D}_{\text{context}}$ to densify the data distribution around $\mathcal{D}_{\text{train}}$.
As the mixing of $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{context}}$ is meant to densify the sparse training data, we only perform the mixing at train time. At test time, the set function $\mu_{\lambda}$ receives an input as a singleton set and performs no mixing with any exogeneous distribution (see Figure 9(a) for train and test-time diagram).
@@ 3 Experiments
Model w/ ctx. w/ bilevel $y^{(\text{mvalid})}$ HIVPROT DPP4 NK1 count vector bit vector count vector bit vector count vector bit vector Ours (Deepsets) rand $1.147\pm 0.055$ $1.012\pm 0.025$ $1.342\pm 0.019$ $1.344\pm 0.039$ $0.45\pm 0.01$ $0.494\pm 0.006$ Ours (Set Trans.) rand $1.193\pm 0.038$ $1.008\pm 0.03$ $1.34\pm 0.026$ $1.359\pm 0.013$ $0.482\pm 0.002$ $0.487\pm 0.009$ Ours (MLP) rand $1.043\pm 0.013$ $1.049\pm 0.013$ $1.42\pm 0.009$ $1.269\pm 0.017$ $0.435\pm 0.002$ $0.466\pm 0.004$ Ours (Deepsets) real $0.506\pm 0.061$ $0.381\pm 0.018$ $\mathbf{0.970\pm 0.012}$ $0.964\pm 0.024$ $0.495\pm 0.037$ $0.395\pm 0.014$ Ours (Set Trans.) real $0.401\pm 0.02$ $0.811\pm 0.111$ $1.123\pm 0.024$ $0.975\pm 0.016$ $0.449\pm 0.021$ $0.401\pm 0.017$ Ours (Deepsets) rand $0.555\pm 0.096$ $\mathbf{0.364\pm 0.018}$ $0.984\pm 0.018$ $\mathbf{0.963\pm 0.017}$ $0.455\pm 0.016$ $\mathbf{0.376\pm 0.008}$ Ours (Set Trans.) rand $\mathbf{0.39\pm 0.011}$ $0.726\pm 0.159$ $1.121\pm 0.037$ $0.986\pm 0.021$ $\mathbf{0.429\pm 0.01}$ $0.397\pm 0.015$
@@
In addition to the baselines presented in (Klarner et al., 2023), we further implement and evaluate two related state-of-the-art interpolation methods—Mixup (Zhang et al., 2017) and Manifold Mixup (Verma et al., 2019)—to empirically demonstrate the capability of our set function $\mu_{\lambda}$. As mixup performs the mixing procedure in the input space, there are no parameters to train for bilevel optimization. We therefore report two variants of mixup (plain and outlier exposure) where outlier exposure is exposed to the same $\mathcal{D}_{\text{mvalid}}$ with random labels as described above. Manifold mixup performs the mixup procedure in the feature space between layers, therefore we can straightforwardly apply our bilevel optimization procedure to manifold mixup. Therefore, we report two variants of manifold mixup (plain and bilevel optimization).
@@ 4 Analysis
To visually assess how our model leverages $\mathcal{D}_{\text{context}}$ to densify $\mathcal{D}_{\text{train}}$, and to compare model behaviors on out-of-distribution (OOD) inputs, we visualize the penultimate-layer embeddings $\mathbf{z}_{i}=x_{i}^{(L-1)}$ , using t-SNE (Van der Maaten & Hinton, 2008) across three models that interpolate $\mathcal{D}_{\text{train}}$ with $\mathcal{D}_{\text{context}}$: Ours, Mixup (w/ outlier exposure), and Manifold Mixup (w/ bilevel optim.), all trained on the DPP4 bit-vector dataset. We randomly sample $x_{i},y_{i}\sim\mathcal{D}_{\text{train}}$, $C_{i}=\{c_{ij}\}_{j=1}^{100}\sim\mathcal{D}_{\text{context}}$, $x_{i}^{\text{ood}}\sim\mathcal{D}_{\text{ood}}$, where $\mathcal{D}_{\text{ood}}$ signifies an out-of-distribution dataset which was never seen during training. In Figure 2, each color indicates a penultimate-layer embedding $\mathbf{z}\in\mathbf{Z}$ under different input configurations. We denote the embeddings as follows: $\mathbf{Z}_{\text{joint}}$ for orange ($\{x_{i},C_{i}\}$), $\mathbf{Z}_{\text{input}}$ for blue ($x_{i}$ alone), $\mathbf{Z}_{\text{context}}$ for green ($C_{i}$ alone), and $\mathbf{Z}_{\text{ood}}$ for purple (OOD samples $x_{i}^{\text{ood}}$). Additional t-SNE results and details are in Appendix B.